40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15990     Entry time: Fri Apr 2 01:26:41 2021
Author: gautam 
Type: Update 
Category: Electronics 
Subject: REFL55 chain checkout again, seems fine 

[koji, gautam]

Summary:

We could not find problems with any individual piece of the REFL55 electronics chain, from photodiode to ADC.  Nevertheless, the PRMI fringes witnessed by REFL55 is ~x10 higher than ~two weeks ago, when the PRMI could be repeatably and reliably locked using REFL55 signals (ETMs misaligned).

Details:

  1. Koji prepared a spare whitening board. However, before he swapped it in, he checked the existing board and found no problems with it.
    • 20mV input signal, 100 Hz was injected into the two REFL55 channels on the whitening board.
    • The flat whitening gain was set to +45 dB.
    • The signal levels seen in CDS was consistent with what is expected given an ADC conversion factor of 3276.8 cts/V.
  2. Tried putting the REFL55 demodulated outputs into the next two channels, 5&6, (currently unused) on the same whitening board.
    • After setting the whitening gains of these two channels also to +18dB, the saturation of the ADCs when the PRMI was fringing persisted.
  3. With the dark noise of the whitening filter, we enabled/disabled the on board frequency dependent whitening, and reasoned that the time domain increase in RMS seemed reasonable. So we decided to investigate parts of the electronics chain upstream of the whitening board, since we couldn't find anything obviously wrong with the whitening board.
  4. Injected -10dBm RF signal (=0.2 Vpp) into the RF input on the REFL55 demod board, and saw ~3500 cts-pp signal in CDS. This is totally consistent with my recent characterization of 16,000 cts/V for this demod board at the "nominal" + 18dB whitening gain setting. So the demodulator seems to function as advertised.
  5. Decided to repeat my test of using the Jenne laser to test the whole chain end-to-end.
    • In summary, we recovered the results (RF transimpedance of the PD, and signal levels in CDS for a known AM determined by the reference NF1611 PD) I reported there.
    • So it would seem that the entire REFL55 electronics chain performs as expected.
    • The only remaining explanation is that the optical gain of the PRMI has increased - but how?? 
    • Similar jumps in the REFL55 signal levels have occurred multiple times in the past, and each time, I was able to recover the "nominal" performance by this procedure (though I have no idea why that should work at all).
    • So I am highly skeptical that this has anything to do with the IFO optical gain, but that is the only difference between our AM laser based test and the "live" operating conditions when the signals are saturated.

Discussion and next steps:

Q: Koji asked me what is the problem with this apparent increased optical gain - can't we just compensate by decreasing the whitening gain?
A: I am unable to transition control of the PRMI (no ETMs) from 3f to 1f, even after reducing the whitening gain on the REFL55 channels to prevent the saturation. So I think we need to get to the bottom of whatever the problem is here.

Q: Why do we need to transfer the control of the vertex to the 1f signals at all?
A: I haven't got a plot in the elog, but from when I had the PRFPMI locked last year, the DARM noise between 100-1kHz had high coherence with the MICH control signal. I tried some feedforward to try and cancel it but never got anywhere. It isn't a quantitative statement but the 1f signals are expected to be cleaner?

Koji pointed out that the MICH signal is visible in the REFL55 channels even when the PRM is misaligned, so I'm gonna look back at the trend data to see if I can identify when this apparent increase in the signal levels occurred and if I can identify some event in the lab that caused it. We also discussed using the ratio of MICH signals in REFL and AS to better estimate the losses in the REFL path - the Faraday losses in particular are a total unknown, but in the AS path, there is less uncertainty since we know the SRM transmission quite precisely, and I guess the 6 output steering mirrors can be assumed to be R=99%. 

ELOG V3.1.3-